# Scikit-fingerprints: easy and efficient computation of molecular fingerprints in Python

### v3

@@ 2.2.5 Code quality and CI/CD
Any changes to the documentation are automatically deployed to the GitHub Pages. New package versions are deployed to PyPI by using GitHub Releases, with new changes description. Internally, this uses a GitHub Actions workflow and creates a Git tag on the commit used in the given release. scikit-fingerprints can be installed via pip by running pip install scikit-fingerprints.
@@ 3.3 Molecular property prediction
Inputs can be any sequences that consist of SMILES strings or RDKit Mol objects, e.g. Python lists or Pandas series. Since ECFPFingerprint is a stateless transformer class, it uses an empty .fit() method in the pipeline. The code is also parallelized, requiring only the n_jobs parameter.

### v4

@@ 2.2.5 Code quality and CI/CD
Any changes to the documentation are automatically deployed to the GitHub Pages. New package versions are deployed to PyPI by using GitHub Releases, with new changes description. Internally, this uses a GitHub Actions workflow and creates a Git tag on the commit used in the given release. scikit-fingerprints can be installed with pip by running pip install scikit-fingerprints.
@@ 3.3 Molecular property prediction
Inputs can be any sequences that consist of SMILES strings or RDKit Mol objects, e.g. Python lists or Pandas series. As ECFPFingerprint is a stateless transformer class, it uses an empty .fit() method in the pipeline. The code is also parallelized, requiring only the n_jobs parameter.
@@ Acknowledgements
Appendix: scikit-mol comparison
This is an additional section, added post-publication and not contained in the published journal version.
Here, we also compare to the scikit-mol library [96], summarized in Table 5.
scikit-mol scikit-fingerprints
Language Python Python
pip-installable Yes Yes
Last PyPI update 2023 2024
Number of fingerprints 8 31
scikit-learn compatible Yes Yes
Parallelism Yes Yes
Pre-commit hooks No Yes
Code quality tools No Yes
Security tools No Yes
Integrated datasets No Yes
Easy proprietary usage
Yes
(LGPL-3.0)
Yes
Table 5: Comparison of scikit-fingerprints with scikit-mol.
Both libraries are written in Python, are installable with pip, and offer a scikit-learn compatible API. In terms of licensing, both enable easy proprietary usage.
While both libraries parallelize fingerprints computation, scikit-mol relies on the built-in Python multiprocessing, which serializes all objects with Pickle when passing them between processes. In scikit-fingerprints, we use the Joblib implementation, which optimizes the serialization and deserialization of NumPy arrays. They are the main result of fingerprint calculation, and the largest objects passed between processes. Hence, their optimization may result in efficiency gains, particularly for large datasets. Joblib also allows seamless switching to distributed computation, since it supports Dask clusters as computation backend.
scikit-fingerprints has much larger implementation scope, with almost four times the number of fingerprints in scikit-mol, and integration of commonly used datasets. Code quality tools, pre-commit hooks, and security tools are not used in scikit-mol, whereas scikit-fingerprints includes them in the CI/CD process.